OpenStudy (anonymous):

Please help! Subtract and simplify completely. 1/3y-2/4x

OpenStudy (saifoo.khan):

Lets Help!

OpenStudy (saifoo.khan):

wht will be the LCM?

OpenStudy (saifoo.khan):

can u guess it?

OpenStudy (anonymous):

\[\frac{1}{3y}-\frac{2}{4x}=\frac{1}{3y}-\frac{1}{2x}=\frac{2x-3y}{6xy}\]